OpenStudy (anonymous):

Which of the following could be the ratio of the length of the longer leg of a 30-60-90 triangle to the length of its hypotenuse? all that apply. 3:sqrt2, 2sqrt3:4, sqrt3:sqrt3, sqrt2:sqrt3, 2:sqrt2

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

which is the same as \[2\sqrt{3}:4\]

OpenStudy (anonymous):

wouldnt 1:sqrt3 be an answer also?